OpenStudy (anonymous):

5 (h - (-1)) divided by 2 --------------------- 7z 5 (h - (-1)) divided by 2 --------------------- 7z @Mathematics

OpenStudy (saifoo.khan):

\[\frac{5(h-(-1)) \div 2}{7z}\]Like that?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

Using PEDMAS. \[\frac{5(h+1)\div2}{7z}\] \[\frac{5h+5 }{7z} \times \frac12\] \[\frac{5h+5}{14z}\]

OpenStudy (anonymous):

Thanks Again

OpenStudy (saifoo.khan):

uw again.